OpenStudy (anonymous):

write an eqaution in slope-intercept form of the line that passes through (4,8) and (0,-2) a.y=5/2x-2 b.y=-5/2x+8 c.y=2/5x-6 d.y=-2/5x+4

OpenStudy (ash2326):

equation of a line passing through points (x2,y2) and (x1,y1) (y-y2)/(x-x2)=(y2-y1)/(x2-x1) here the two points are 4,8 and 0,-2 (y-8)/(x-4)=(8-(-2))/(4-0) (y-8)/(x-4)=2.5 y-8=2.5x -10 y=2.5x-2 slope is -2.5 and y-intercept is -2 or y=5/2x-2 option A